OpenStudy (anonymous):

A function has the rule "multiply by two, and then subtract one." What is the output when the input is 7? 4 twelve thirteen fifteen

OpenStudy (anonymous):

x \(\times\) 2 - 1 x = 7 7 \(\times\) 2 - 1 Can you solve that @housekill11

OpenStudy (anonymous):

c??

OpenStudy (anonymous):

no wait B

OpenStudy (anonymous):

13 so c

OpenStudy (anonymous):

am i right ?

OpenStudy (anonymous):

Yes, C. @housekill11